7. A ball is thrown vertically upward with a speed of 7.1 m/s. Find:



(9 points)



a. Maximum height reached the ball



b. Its time of flight



C. Velocity the ball would return to its starting point

1
Expert's answer
2022-05-05T13:54:12-0400

Given:

"v_0=7.1\:\rm m/s"

"g=9.8\:\rm m/s^2"


(a)

"h_{\max}=\frac{v_0^2}{2g}=\frac{7.1^2}{2*9.8}=2.6\:\rm m"

(b)

"t=2*\frac{v_0}{g}=2*\frac{7.1}{9.8}=1.4\:\rm s"

(c)

"v_f=v_0=7.1\:\rm m/s"